Find the number of joules needed to heat 37grams of water 78 degees C.
1
Expert's answer
2015-09-10T10:07:49-0400
To heat 1 g on 1 degree C we need 1 cal.
So to heat 37 g on 78 degrees C we need:
C=37*78=2886 cal.
1 cal.=4.184 J
So, 2886 cal.=2886*4.184=12075.024 J.